Grade 12Physical Chemistry

if 10 ml of 10 ^-6 M HCI solution is mixed with 90 ml H2O pH will change approximately by

Darlene Nazareth
10 Years agoGrade 12

2 Answers

Bhavya
10 Years ago
In order to calculate pH, we must first know the resulting molarity of the mixture.
By applying M1V1 + M2V2 = MfVf
10*10-6 + 90*10-7 = M*(90+10) = M*100
10-7 (1+90) = M*100
10-7 (91/100) = M
0.91 *10-7 = M
pH = -log[H+] = – log[0.91*10-7]
= 7.04
Aditya agrahari
6 Years ago
First we calculate the resulting  molarity of solution, 
Using M1V1+ M2V2=M3V3, Here final volume V3 = 100ml .
On calculating we get the value M3 to be 19*10^-8M
pH= -log[H+] = - log[19*10^-8] = 7.72
therefore pH is changed by 0.7 approximately